A rock is thrown upward from the level ground in such a way that the maximum height of its flight is equal to its horizontal range R.In terms of its original range R, what is the range Rmax the rock can attain if it is launched at the same speed but at the optimal angle for maximum range?


1
Expert's answer
2021-08-19T14:23:22-0400

"R=\dfrac{u^2\sin{2\theta}}{g}\\
H=\dfrac{u^2\sin^2{\theta}}{2g}\\
\dfrac{u^2\sin{2\theta}}{g}=\dfrac{u^2\sin^2{\theta}}{2g}\\
2\sin{2\theta}=\sin^2{\theta}\\
4\sin{\theta}\cos{\theta}=\sin^2{\theta}\\
\tan{\theta}=4\\
\theta=\arctan{4}=76\degree\\
R=\dfrac{u^2\sin{2\theta}}{g}\\
At\ same\ speed\ and\ optimal\ angle\ for\\ maximum\ range\\
\theta=45\degree\\
R_{max}=\dfrac{u^2}{g}\\
R=R_{max}\sin{2\theta}\\
R_{max}=\dfrac{R}{\sin{2\theta}}"